OpenStudy (anonymous):

Consider the so called general linear group GL(n, R) = {A ∈ Rn×n: det A is not equal 0} with the matrix multiplication as the group operation. Prove that the subset SL(n, R) = {A ∈ Rn×n: det A = 1} is a normal subgroup of GL(n, R).

OpenStudy (anonymous):

We at least know it is a subset, given that \(\det A = 1 \implies \det A \neq 0\).

OpenStudy (anonymous):

Given that \(\det(AB) = \det(A)\det(B)\), we also know it'll be closed.

OpenStudy (anonymous):

I don't remember all the properties of a normal subgroup.

OpenStudy (anonymous):

H is normal of G if xHx^-1=H

OpenStudy (rsadhvika):

You just need to show \(\det(PAP^{-1} )= 1\)

OpenStudy (rsadhvika):

where \(P \in GL(n,R)\) and \(A \in SL(n,R)\)

OpenStudy (anonymous):

how can we do that? can you help me please?

OpenStudy (rsadhvika):

use below : \(\det(P^{-1}) = \dfrac{1}{\det(P)}\)

OpenStudy (anonymous):

is it the same as DET B.DET A. DET P-1=1

OpenStudy (rsadhvika):

Yes \[\det(PAP^{-1} )\\~\\= \det(P)\det(A)\det(P^{-1})\\~\\ = \det(P)\det(A)\dfrac{1}{\det(P)} \\~\\= \det(A) \\~\\=1\]